ok!!! i have an exam coming up and i need help in geometry!!! How would i find using the pthagorean theam values of x in a triangle with a 90 degree angle and side a=6 side b=x and side c=8?
Answer by stanbon(75887)

values of x in a triangle with a 90 degree angle and side a=6 side b=x and side c=8?
Pythagoras: c^2 = a^2 + b^2 where c is the hypotenuse of a rt. triangle.
In your problem c=8, a=6, b=x
8^2= 6^2 + x^2
x^2= 64-36
x^2= 28
x=2sqrt7
Cheers,
Stan H.